# Designing Elegant Bell Inequalities

### v2

@@ Abstract
Elegant Bell inequality is well known for its much exploited property, being maximally violated by maximal entanglement, mutually unbiased bases, and symmetric informationally complete positive operator-valued measure elements. It is the only one with such property known so far. We present a method to construct Bell inequalities with violation feature analogous to original elegant Bell inequality in high dimension from a simple analytic quantum bound. A Bell inequality with such feature is derived in three dimension for the first time. It shows larger violation than existing Bell inequalities of similar classes while requiring arguably small number of measurements.
I Introduction
Performing local measurements on the shared system, distant subsystems can reveal a quantum correlation that cannot be reproduced by any local hidden variable theoretical explanation. Such nonlocality of the correlation can be tested via the violation of Bell inequalities [1, 2]. What Bell inequality can tell us about the system goes beyond the scope of the mere existence of nonlocal correlation. It has been investigated the method to exploit Bell’s theorem for further certifying the properties of given system in so-called device-independent (DI) approach [3, 4]. This approach leads to various applications, including DI quantum key distribution [4, 5, 6, 7, 8] and DI certification of randomness [9, 10, 11].
Reflecting on the aforementioned possibilities provided by Bell inequalities, it is of prime importance to derive them such that their maximal violations are obtained from desirable features, for example maximal entanglement [12, 13, 14, 15, 16, 17, 18, 19, 20], mutually unbiased bases (MUBs) [19, 20], or symmetric informationally complete POVM elements (SICs) [21, 20]. As is well known, Clauser-Horne-Shimony-Holt Bell inequality is the case in which the maximal violation can be obtained from maximal entanglement and Pauli operators having MUBs as their eigenstates. There exists another bipartite Bell inequality so-called Gisin’s elegant Bell inequality (EBI) [21],
@@ I Introduction
also optimally violated by maximally entangled state. The expectation value of the product outcomes of Alice’s $x$-th measurement and Bob’s $y$-th measurement is denoted by $E_{xy}$. Quantum bound for EBI is proven as $64\sqrt{3}\simeq 6.9282>$ 6 [22]. What makes EBI more interesting is that its violation is obtained from the case where the local measurements of each subsystem is constructed by MUBs and SICs respectively. As first remarked by Gisin, the geometrical symmetries in the optimal measurement settings of EBI are clearly revealed when they are mapped on Bloch’s sphere [21]. Three eigenstates defining three projective measurements of Alice form three mutually orthogonal vectors on Bloch sphere and four eigenstates of Bob’s side make a tetrahedron [21]. Original EBI has been used in several fields of quantum information, including DI randomness certification [22, 23], DI certification of non-projective measurement [24, 25], and for other tasks [26, 27, 28]. Self-testing property of EBI has also been studied in [29].
MUBs and SICs are important resources in quantum cryptography [30, 31, 32, 22, 23]. There has been efforts to derive Bell inequalities for MUBs or SICs exploiting the advantage of high dimension [19, 20]. For DI randomness certification, for which high dimension provides natural advantage [22], Bell inequalities for SICs has been used [22, 23, 20]. EBI is distinguished from ones either for MUBs or SICs as it is optimally violated by both of them.
We address the problem of deriving Bell inequalities with violation feature analogous to original EBI in high dimensions. The problem is closely related to the question of what benefit for Bell nonlocality would come from simultaneous consideration of the symmetries of MUBs and SICs. It is worth noting that although EBI is originally derived in the different context [33, 21], we focus on the generalization of the property, being maximally violated by maximal entanglement, MUBs and SICs. Such generalization of EBI has not much been investigated so far to our knowledge. We approach the raised problem by defining a generalized type of correlation as follows.
Definition. For prime $d\geq 3$, generalized EBI correlation is defined as the bipartite correlation obtained from the quantum realization: (i) maximally entangled state $\left|\phi_{d}^{+}\right>:=(1/\sqrt{d})\sum_{\alpha=0}^{d-1}\left|\alpha\alpha\right>$, (ii) Alice’s $d^{2}-1$ observables comprising $\mathcal{W}_{d}\setminus\mathds{1}$ and (iii) Bob’s $d^{2}$ observables having $d^{2}$ SICs in their eigenspaces.
In condition (ii), $\mathcal{W}_{d}:=\{X^{m}Z^{n}|m,n\in[0,d)\}$ is the set of $d^{2}$ Weyl-Heisenberg (WH) operators defined with $X:=\sum_{\alpha=0}^{d-1}\left|\alpha+1\right>\left<\alpha\right|$, $Z:=\sum_{\alpha=0}^{d-1}\omega^{\alpha}\left|\alpha\right>\left<\alpha\right|$ and $\omega:=e^{2\pi i/d}$. The integer domain $\{0,1,\ldots,d-1\}$ is denoted by $[0,d)$ and similar abbreviation is to be used throughout this work. Note that $X^{0}Z^{0}=\mathds{1}$ is not included in the condition (ii). Conditions (i) and (iii) are natural generalization of optimal quantum realization of original EBI. As for the generalization of Alice’s optimal observables, two interpretations are possible. The set of Alice’s optimal observables for EBI, $\sigma_{x},\sigma_{y},\sigma_{z}$, [22] can be viewed either as “$d+1$ observables” defining $d+1$ sets of MUBs, or as “$d^{2}-1$ unitary operator bases” [34] for a traceless operator $\vec{a}\cdot\vec{\sigma}$ defined with a three-dimensional unit vector $\vec{a}$ and $\vec{\sigma}:=(\sigma_{x},\sigma_{y},\sigma_{z})$ when $d=2$. The above duality arises as $d+1=d^{2}-1$ when $d=2$. Condition (ii) clearly generalizes the latter. Under condition (ii), complete MUBs are realized in eigenspace of $d+1$ among $d^{2}-1$ observables of Alice, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [35]. Moreover, the eigenbases of all $d^{2}-1$ observables of Alice comprises the $d-1$ same sets of complete $d+1$ MUBs. It can be easily seen from the fact that $XZ^{n}$ and $X^{m}Z^{mn}$ has the same eigenbases for all $m$ and $\mathcal{W}=\{X^{m}Z^{mn}|m,n\in[0,d)\}$ for prime $d$. In the following, we will further explain the reason for the choice of optimal observables of Alice as unitary operator bases.
Figure 1: On the shared system, two subsystems, Alice and Bob, respectively performs one of $d^{2}-1$ and $d^{2}$ measurements outputting one of $d$ outcomes.
Another insight that underlies our approach is to exploit $systematic$ and $versatile$ features of WH group covariant SICs. For brevity, consider the parameterization $W_{dp+q}:=X^{p}Z^{q}$, such that $\mathcal{W}_{d}=\{W_{j}|j\in[0,d^{2})\}$. Group covariant SICs is derived by acting group elements on a single state, so-called fiducial vector [36]. For the case of WH group covariant SICs, group elements $W_{j}$ generates $d^{2}$ SICs on the support of fiducial state $\left|\varphi\right>$, $\{W_{j}\left|\varphi\right>|j\in[0,d^{2})\}$. Weyl-Heisenberg group covariance plays crucial role for satisfying the condition (iii) and provides systematic feature for our formalism. Versatility of our method is from the fact that it can be applied only if WH group covariant SICs for given local dimension exists. The existence of SICs are known in all dimension up to $d=193$ [37]. More importantly, all group covariant SICs is covariant to WH group in prime dimensions [38].
II Preliminaries
@@
We consider the measurement scenario of two subsystems Alice and Bob respectively performing one of possible $d^{2}-1$ and $d^{2}$ number of $d$ outcome measurements $A_{x}$ and $B_{y}$ on the shared system. Measurements are labelled with $x\in\{1,2,\ldots,d^{2}-1\}$ and $y\in\{0,1,\ldots,d^{2}-1\}$. Outcomes of $A_{x}$ and $B_{y}$ are respectively distinguished with indices $\alpha,\beta\in\{0,1,\ldots,d-1\}$. The scenario is illustrated in Fig. 1. Probability of obtaining the outcomes $\alpha$ and $\beta$ respectively from settings $x$ and $y$ is denoted by $P(\alpha\beta|xy)$. A statistics of the Bell test can be described with correlation, $\mathbf{p}:=\{P(\alpha\beta|xy)|\forall\alpha,\beta,x,y\}$. Quantum correlation allows
@@
where $\rho_{AB}$ is the bipartite density matrix of Alice and Bob. Measurement operators, $\{G_{\alpha}^{x}\}$ and $\{H_{\beta}^{y}\}$, define the measurements of Alice and Bob respectively. Bell inequality is used to show that there is quantum correlation which does not allow the local hidden variable (LHV) theoretical description of the correlation [1]. Geometrically, LHV correlation region forms a polytope, $\mathcal{L}$, and quantum region, $\mathcal{Q}$, is known to include it [2].
The generic form of so-called Bell expression [2] without marginal probabilities can be defined as,
@@
where $n\in(1,d)$ and $g_{x,y}^{\alpha,\beta}\in\mathbb{R}$ is the real weight for each probability. For simplicity, the summations over all possible values are to be denoted without limits. The correlator $\langle A_{x}^{n}B_{y}^{n}\rangle=\sum_{\alpha,\beta=0}^{d-1}\omega^{n(\alpha+\beta)}P(\alpha\beta|xy)$, where the outcomes of the measurements $A_{x}$ and $B_{y}$ are set as $\omega^{\alpha},\omega^{\beta}\in\{1,\omega,\omega^{2},\ldots,\omega^{d-1}\}$. And the coefficient $g_{x,y}^{\alpha,\beta}:=\sum_{n}f_{x,y}^{n}\omega^{n(\alpha+\beta)}$ The condition $f_{x,y}^{d-n}:=(f_{x,y}^{n})^{*}$ guarantees the real-valuedness of $S$. A expression is specified by a $(d^{2}-1)\times d^{2}$ coefficient matrices, $F_{n}$, whose entries are $\{f_{x,y}^{n}\}$. We define $x$-th row and $y$-th column of $F_{n}$ respectively as $\mathbf{r}_{x}^{n}:=(f_{x,0}^{n},f_{x,1}^{n},\ldots,f_{x,d^{2}-1}^{n})^{\intercal}$ and $\mathbf{c}_{y}^{n}:=(f_{1,y}^{n},f_{2,y}^{n},\ldots,f_{d^{2}-1,y}^{n})^{\intercal}$ such that $F_{n}=[\mathbf{r}_{1}^{n},\mathbf{r}_{2}^{n},\ldots,\mathbf{r}_{d^{2}-1}^{n}]^{\intercal}=[\mathbf{c}_{0}^{n},\mathbf{c}_{1}^{n},\ldots,\mathbf{c}_{d^{2}-1}^{n}]$. In terms of $F_{n}$, real-valuedness condition is re-written as $F_{n}=F_{d-n}^{*}$.
@@
III Main result

### v3

# Designing generalized elegant Bell inequalities in high dimension from a quantum bound

@@ Abstract
Elegant Bell inequality is well known for its distinctive property, being maximally violated by maximal entanglement, mutually unbiased bases, and symmetric informationally complete positive operator-valued measure elements. Despite its significance in quantum information theory demonstrated based on its unique violation feature, it remains the only known one with the characteristic. We present a method to construct Bell inequalities with violation feature analogous to elegant Bell inequality in higher local dimension from a simple analytic quantum bound. A Bell inequality with the generalized violation feature is derived in three dimension for the first time. It exhibits larger violation than existing Bell inequalities of similar classes, including the original elegant Bell inequality, while requiring arguably small number of measurements.
1 Introduction
Performing local measurements on the shared system, distant subsystems can reveal a quantum correlation that cannot be reproduced by any local hidden variable theoretical explanation. Such nonlocality of the correlation can be tested via the violation of Bell inequalities [1, 2]. The implications of Bell inequalities extend beyond merely confirming the existence of nonlocal correlations. Researchers have investigated methods to exploit Bell’s theorem for certifying the properties of given system through the device-independent (DI) approach [3, 4, 5]. This approach leads to various applications, including DI quantum key distribution [4, 6, 7, 8, 9] and DI certification of randomness [10, 11, 12].
Reflecting on the aforementioned possibilities provided by Bell inequalities, it is of prime importance to derive them such that their maximal violations are obtained from desirable features, for example maximal entanglement [13, 14, 15, 16, 17, 18, 19, 20, 21], mutually unbiased bases (MUBs) [20, 21], or symmetric informationally complete quantum measurements (SICs) [22, 21].
As is well known, Clauser-Horne-Shimony-Holt Bell inequality is the case in which the maximal violation can be obtained from maximal entanglement and Pauli operators having MUBs as their eigenstates. There exists another bipartite Bell inequality so-called Gisin’s elegant Bell inequality (EBI) [22],
@@ 1 Introduction
also optimally violated by maximally entangled state. The expectation value of the product outcomes of Alice’s $x$-th measurement and Bob’s $y$-th measurement is denoted by $E_{xy}$. Quantum bound for EBI is proven as $64\sqrt{3}\simeq 6.9282>$ 6 [23]. What makes EBI more interesting is that its violation is obtained from the case where the local measurements of each subsystem is constructed by MUBs and SICs respectively. As first remarked by Gisin, the geometrical symmetries in the optimal measurements of EBI are clearly revealed when they are mapped on Bloch’s sphere [22]. Three eigenstates defining three projective measurements of Alice form three mutually orthogonal vectors on Bloch sphere and Bob’s four eigenstates make a tetrahedron [22]. EBI has been applied for several tasks in quantum information field, including DI randomness certification [23, 24], DI certification of non-projective measurement [25, 26], and other tasks [27, 28, 29]. Self-testing property of EBI is studied in [30].
MUBs and SICs are not only relevant for the formulation of quantum mechanics [31], but also for the extensive quantum information processing tasks [23, 24, 32, 33, 34, 35, 36, 37, 38, 39, 40]. There has been efforts to derive Bell inequalities for MUBs or SICs exploiting the high dimensional system [20, 21]. For DI randomness certification, for which high dimension provides natural advantage [23], Bell inequalities for SICs has been used [21, 23, 24]. EBI is distinguished from those for either MUBs or SICs as it is optimally violated by both. Randomness certification is a notable application of EBI. DI certification scheme of optimal randomness from one entanglement bit is suggested based on EBI [23, 24]. Considering the above, investigating the desirable feature of Bell inequalities analogous to EBI in high dimension is expected to be intriguing from the application perspective beyond its implications just for extending the mathematical symmetries.
We address the problem of deriving Bell inequalities in higher dimensions with violation features analogous to EBI. This investigation is closely related to understanding how the simultaneous consideration of MUB and SIC symmetries could enhance Bell nonlocality. Although EBI originated in a different context [22, 41], our focus lies in generalizing its distinctive property of achieving maximal violation through maximal entanglement, MUBs, and SICs. Such generalization of EBI remains largely unexplored in the literature.
This article is structured as follows. In Sec. 2, we briefly introduce MUBs and SICs, key concepts of our result. Bell scenario and notations defining it are given in the same section. In Sec. 3, the generalized violation feature inspired from EBI is defined. It is denoted under a term, generalized EBI correlation. In Sec. 4, we present the main result deriving the form of correlation function maximized with the generalized EBI correlation. The corresponding quantum bound is derived in an analytic form. In Sec. 5, to show the applicability of our result, we derive Gisin’s EBI and a novel Bell inequality for three dimensional generalized EBI correlation. We evaluate the violation of the Bell inequality in comparison with established results.
2 Preliminaries
2.1 MUBs and SICs
We briefly introduce the definition of MUBs and SICs, along with their properties highly relevant to this work. Comprehensive reviews of MUBs and SICs, including their fundamental concepts and properties, can be found in [35, 42].
2.1.1 MUBs
Normalized bases $\{\left|a_{i}\right>\}_{i=0}^{d-1}$ and $\{\left|b_{j}\right>\}_{j=0}^{d-1}$ are said to be mutually unbiased when they satify
|$a_{i}|b_{j}\right>\right|^{2}=\frac{1}{d}$ (1)
for all combinations of $i$ and $j$. Problem of finding MUBs in all Hilbert space dimesion is still open. However, it is known that whenever the prime power dimension $d$ is considered there are at most $d+1$ MUBs [40], referred to as the complete set.
The complete set of MUBs in prime $d$ dimension can be constructed using Weyl-Heisenberg (WH) group elements. The shift $X:=\sum_{\alpha=0}^{d-1}\left|\alpha+1\right>\left<\alpha\right|$ and the phase $Z:=\sum_{\alpha=0}^{d-1}\omega^{\alpha}\left|\alpha\right>\left<\alpha\right|$ operators generate the WH group $\{\omega^{n}X^{p}Z^{q}|n,p,q\in[0,d)\}$ with $\omega:=e^{2\pi i/d}$. The integer domain $\{0,1,\ldots,d-1\}$ is denoted by $[0,d)$ and similar abbreviation is to be used throughout this work. Note that $X^{d}=Z^{d}=\mathds{1}$ and the canonical commutation relation (CCR) holds, i.e. $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$. A subset of WH group without phase factor also forms a group,
$\mathcal{W}(d)=\{X^{m}Z^{n}|m,n\in[0,d)\}.$ (2)
The complete MUBs can be obtained as the eigenvectors of $d+1$ elements in $\mathcal{W}(d)$, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [43].
2.1.2 SICs
Suppose a set of $d^{2}$ normalized states $\{\left|e_{i}\right>\}_{i=0}^{d^{2}-1}$ satisfying the relation
|$e_{i}|e_{j}\right>\right|^{2}=\frac{1}{d+1}$ (3)
for all $i\neq j$. The above constant overlap relations reflects the geometrical symmetry among $d^{2}$ states. They are equiangular lines in $\mathds{C}^{d}$ [42]. The states define a symmetric informationally complete positive operator-valued measure (POVM), $\left\{\frac{1}{d}\left|e_{i}\right>\left<e_{i}\right||i\in[0,d^{2})\right\}$ with additional condition that all the POVM elements must be summed up to become identity operator. The POVM is informationally complete in the sense that its statistics characterizes the quantum state $\rho$ in complex $d$ dimension upon which it is carried on. The $d^{2}$ number of elements are necessary for a informationally complete POVM. In this article, a SICs specifically denotes the $d^{2}$ unit vectors $\{\left|e_{i}\right>\}_{i=0}^{d^{2}-1}$ satisfying the relation (3).
It is interesting that SICs are also closely related to WH group. WH operator bases hence provide a suitable common ground for the analysis of MUBs and SICs in high dimension. It is the very reason why we adopt the WH bases. SIC-POVM is group covariant if there is a group $\{U_{i}\}$ having unitary operators as its elements and SICs is generated from operating the elements upon a single vector (or state), so-called fiducial vector $\left|f\right>$, i.e. when $\{U_{i}\left|f\right>\}$ is SICs [45]. The group covariance of SICs plays a significant role in the problem of SICs generation [42].
2.2 Bell Scenario
@@
We consider the measurement scenario of two subsystems Alice (Bob) performing one of possible $d^{2}-1$ ($d^{2}$) number of $d$ outcome measurements $A_{x}$ ($B_{y}$) on their shared system. Measurements are labelled with $x\in\{1,2,\ldots,d^{2}-1\}$ and $y\in\{0,1,\ldots,d^{2}-1\}$. Outcomes of $A_{x}$ and $B_{y}$ are respectively distinguished by indices $\alpha\in\{0,1,\ldots,d-1\}$ and $\beta$ in the same domain. The scenario is illustrated in Fig. 1. Probability of obtaining the outcomes $\alpha$ and $\beta$ respectively from settings $x$ and $y$ is denoted by $p(\alpha\beta|xy)$. A statistics of the Bell test can be described with $\mathbf{p}:=\{p(\alpha\beta|xy)\}_{\alpha,\beta,x,y}$, a correlation. Quantum correlation allows
@@
where $\rho_{AB}$ is a bipartite density matrix of Alice and Bob. The operators, $\{G_{\alpha}^{x}\}$ and $\{H_{\beta}^{y}\}$, define the measurements of Alice and Bob respectively. Bell inequality is used to show that there is quantum correlation, which does not allow the local hidden variable (LHV) theoretical description of the correlation [1]. Geometrically, LHV correlation region forms a polytope, $\mathcal{L}$, and quantum region, $\mathcal{Q}$, includes it [2].
The generic form of so-called Bell expression [2] without marginal terms can be defined as
@@
where $n\in[1,d)$ and $g_{x,y}^{\alpha,\beta}\in\mathbb{R}$ is the weight for each probability. For simplicity, the summations over all possible values are to be denoted without limits. The correlator $\langle A_{x}^{n}B_{y}^{n}\rangle=\sum_{\alpha,\beta=0}^{d-1}\omega^{n(\alpha+\beta)}P(\alpha\beta|xy)$, where the outcomes of the measurements $A_{x}$ and $B_{y}$ are set as $\omega^{\alpha},\omega^{\beta}\in\{1,\omega,\omega^{2},\ldots,\omega^{d-1}\}$. Equivalence of the expressions (4) and (5) are given by the relation $g_{x,y}^{\alpha,\beta}:=\sum_{n}f_{x,y}^{n}\omega^{n(\alpha+\beta)}$ The condition $f_{x,y}^{d-n}:=(f_{x,y}^{n})^{*}$ guarantees the real-valuedness of the expression. A possible expression is specified by a $(d^{2}-1)\times d^{2}$ coefficient matrices, $F_{n}$, whose entry $(F_{n})_{x,y}$ is defined as $f_{x,y}^{n}$. We define $x$-th row and $y$-th column of $F_{n}$ respectively with column vectors $\mathbf{r}_{x}^{n}:=(f_{x,0}^{n},f_{x,1}^{n},\ldots,f_{x,d^{2}-1}^{n})^{\intercal}$ and $\mathbf{c}_{y}^{n}:=(f_{1,y}^{n},f_{2,y}^{n},\ldots,f_{d^{2}-1,y}^{n})^{\intercal}$ such that $F_{n}=[\mathbf{r}_{1}^{n},\mathbf{r}_{2}^{n},\ldots,\mathbf{r}_{d^{2}-1}^{n}]^{\intercal}=[\mathbf{c}_{0}^{n},\mathbf{c}_{1}^{n},\ldots,\mathbf{c}_{d^{2}-1}^{n}]$. The real-valuedness condition can be re-written as $F_{n}=F_{d-n}^{*}$, in terms of coefficient matrices.
@@
[…]
[scale=0.13]fig_scenario.jpg
Figure 1: The Bell scenario considers the $d^{2}-1$ and $d^{2}$ inputs in local systems outputting one of $d$ outcomes.
3 Generalized EBI correlation
In the introduction, the problem of deriving Bell inequalities analogous to EBI in high dimension is raised. We approach the problem by defining an optimal correlation as follows.
Definition. For prime $d\geq 3$, a generalized EBI correlation is defined as the bipartite correlation obtained from the quantum realization:
{adjustwidth}
.5em0pt (i) $\left|\phi_{d}^{+}\right>:=(1/\sqrt{d})\sum_{\alpha=0}^{d-1}\left|\alpha\alpha\right>$
(ii) Alice’s $d^{2}-1$ observables comprising $\mathcal{W}_{d}\setminus\mathds{1}$
(iii) Bob’s $d^{2}$ observables respectively having each one of the $d^{2}$ unit vectors satisfying (3) as an eigenstate.
Conditions (i)-(iii) define a possible generalization of optimal quantum realization of EBI. Condition (i) is a two-qudit maximally entangled state. Condition (iii) guarantees the realization of at least one SICs in the Bob’s eigenspace. These conditions are straightforwardly derived from the optimal realization of EBI. As for the generalization of Alice’s optimal observables, two interpretations are possible. The set of Alice’s optimal observables for EBI, $\sigma_{x},\sigma_{y},\sigma_{z}$, [23] can be viewed either as “$d+1$ observables” defining $d+1$ sets of MUBs or as “$d^{2}-1$ unitary operator bases” [31]. Condition (ii) clearly generalizes the latter. Note that a traceless operator can be expressed as a familiar form $\vec{a}\cdot\vec{\sigma}$ with the unitary operator bases, $\vec{\sigma}:=(\sigma_{x},\sigma_{y},\sigma_{z})$, when $d=2$ and $\vec{a}$ is a three dimensional unit vector. The above duality arises as $d+1=d^{2}-1$ when $d=2$. Excluding trivial case $X^{0}Z^{0}=\mathds{1}$ makes the number of observables $d^{2}-1$ rather than $d^{2}$ in condition (ii). Complete MUBs are realized in eigenspace of $d+1$ among $d^{2}-1$ observables of Alice, namely $Z$, $X$, $XZ$, …, $XZ^{d-1}$ [43], under condition (ii). Moreover, the eigenstates of all $d^{2}-1$ observables of Alice falls into one of eigenstates of $Z$, $X$, $XZ$, …, $XZ^{d-1}$ under condition (ii). In other words, the eigenspaces of $d^{2}-1$ observables are $d-1$ folded eigenspaces of $d+1$ complete MUBs. (See appendix A for the detailed discussion.)
WH operators play a pivotal role in not only deriving Alice’s optimal observable but also for Bob’s ones. The explicit derivation is given in the following section. Here, we briefly convey the advantages brought by exploiting WH group covariant SICs for satisfying condition (iii). For brevity, consider the parameterization $W_{dp+q}:=X^{p}Z^{q}$ such that $\mathcal{W}(d)=\{W_{j}|j\in[0,d^{2})\}$. For example, $W_{5}=XZ^{2}$ when $d=3$. With the parameterization, WH group covariant SICs are $\{W_{j}\left|\varphi\right>|j\in[0,d^{2})\}$ for a fiducial state $\left|\varphi\right>$. Group covariant SICs enable one to satisfy the condition (iii) in a systematic way. The condition is satisfied by assigning each one of $d^{2}$ states, $W_{j}\left|\varphi\right>$, to a different eigenstate of Bob’s $j$-th measurement. Consequently, Bob’s optimal measurements defined in condition (iii) are derived neatly as (9), in the following section. The WH group covariant SICs also provide versatile nature for our framework. It is based on the fact that our method for deriving the quantum bound can always be applied if WH group covariant SICs of given dimension exists. The existence of SICs is known in all dimension up to $d=193$ [44]. More importantly, all group covariant SICs are covariant to WH group in prime dimension [46] that we mainly consider. In light of the above discussion, exploiting WH group covariance of SICs is arguably reasonable to derive and analyze the Bell nonlocality for generalized EBI correlation.
In the following section, we elaborate on how the WH operator bases also offer advantages in the optimization of the quantum correlation.
4 Main result